OpenStudy (tumblewolf):

I need help with a limit defenition problem. Will fan and medal! I did it in derivative form and my teacher wants it the other way but I can't figure it out. F(x)=2x^-6x+2 and x=2 I have it set up as ((2(2+h)^2-6(2)+2) - (2(2^2)+2)))/h And don't know where to go from here. Thank you!!

OpenStudy (dumbcow):

def of derivative is \[\lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h}\]

OpenStudy (tumblewolf):

I know that much

OpenStudy (dumbcow):

whats the other form that teacher wants?

OpenStudy (tumblewolf):

I'm sorry I wasn't clear! She wants us to use that to find the equation of the tanget line Sorry!

OpenStudy (dumbcow):

oh ok

OpenStudy (dumbcow):

equation of tangent line: \[ y - f(x_0) = f'(x_0) ( x - x_0)\] x_0 = 2

OpenStudy (dumbcow):

also i think you are missing some terms in your setup \[\rightarrow \frac{ (2(h+2)^2 - 6(h+2) +2)) - (2*2^2 - 6(2) +2)}{h}\]

OpenStudy (dumbcow):

you have to simplify this by distributing and combining like terms

OpenStudy (tumblewolf):

So do I need to add the h in with the 6?

OpenStudy (dumbcow):

yes .... everywhere there is an "x" , you must replace with "h+2"

OpenStudy (tumblewolf):

Would the 2 be before the H or does it matter?

OpenStudy (dumbcow):

doesn't matter ..... a+b = b+a

OpenStudy (tumblewolf):

\[((12h+4^2-6h+12+2)-(16-12+2))/h\]

OpenStudy (tumblewolf):

Only 2h

OpenStudy (dumbcow):

ok you need to square the (h+2) first before distributing the 2 \[(h+2)^2 = (h+2)(h+2) = h^2 +2h+2h +4\]

OpenStudy (tumblewolf):

Sorry my computer died!

OpenStudy (tumblewolf):

\[(4+2h)(4+h)-12+6h+2)-(16-12+2)\] Would this be right?

OpenStudy (dumbcow):

no , remember to distribute correctly --> -6(h+2) = -6h - 12 , the negative is multiplied by both terms 2(h+2)^2 ----> square first, then distribute the 2 (look at my last post)

OpenStudy (dumbcow):

2(2^2) = 2(4) = 8 ..... not 16

OpenStudy (tumblewolf):

Oh now I see

OpenStudy (tumblewolf):

\[8+2h-12+6h)+2)-8-12+2\]

OpenStudy (dumbcow):

how do you get 8 +2h ?

OpenStudy (dumbcow):

\[2(h+2)^2 = 2(h^2 +4h+4) = 2h^2 +8h+8\]

OpenStudy (tumblewolf):

it's 2(2+h)^2 so 2 squared is four. four times 2 is 8.

OpenStudy (tumblewolf):

I think I'm just stressed right now and not thinking straight

OpenStudy (dumbcow):

yes that is correct for the constant term.... but you have to include the variable "h" when you square it remember FOIL, multiplying binomials (x+2)(x+2)

OpenStudy (dumbcow):

anyway you should have \[\frac{(2h^2 +8h+8 -6h-12+2) - (8-12+2)}{h}\] \[= \frac{2h^2+2h}{h}\] \[= \frac{h(2h+2)}{h}\] \[=2h +2\] Take limit as h->0 (plug in h=0) \[f'(2) = 2\]

OpenStudy (tumblewolf):

So what would the intercept be?

OpenStudy (tumblewolf):

\[y-(-2)=2(x-2)\]